OpenStudy (anonymous):

Solve -(2x + 4) = 18.

OpenStudy (anonymous):

-2x - 4 = 18 -2x = 18 + 4 -2x = 22 x = -11

OpenStudy (anonymous):

-2x - 4 = 18 -2x = 18 + 4 -2x = 22 x = -11